OpenStudy (anonymous):

can someone check this for me? y=-4x^2-8x-14 x=-3

OpenStudy (anonymous):

@Jack_Prism

OpenStudy (jack_prism):

Correct

OpenStudy (anonymous):

I did something wrong before and got -46...XD

OpenStudy (anonymous):

is this right...? -4x(-3x-3)=-36 8x-3=-24 -24-14=-38 -36--38=2? wait...what am I doing wrong?

OpenStudy (anonymous):

how do I get it to -26 this is making no sense...

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@AloneS

alones (alones):

yes???

OpenStudy (anonymous):

do you know how to get -26 on this equation? y=-4x^2-8x-14 x=-3 theres something I'm doing wrong...

alones (alones):

wait so this is your question??

OpenStudy (anonymous):

yep...XD

OpenStudy (anonymous):

I cant figure out what I'm doing wrong on this...

alones (alones):

wait i'm confused where is ..wait can you show me the question xD i don't understan

OpenStudy (anonymous):

y=-4x^2-8x-14 x=-3 this IS it, I just have to solve it

alones (alones):

do you need to find -26

OpenStudy (anonymous):

I have to find a way to make it equal to -26 I keep getting it wrong, I got -46 but it has to be -26

alones (alones):

O_o you just need to show your work?? for this answer

OpenStudy (anonymous):

yes...XD

alones (alones):

exaclty xD haha okayii i appologise i didn't understand it at first time

OpenStudy (anonymous):

I see lol but I keep getting -46 theres something I'm either doing wrong or missing...

alones (alones):

y = -14 + -8x + -4x2 Solving y = -14 + -8x + -4x2 Solving for variable 'y'. Move all terms containing y to the left, all other terms to the right. Simplifying y = -14 + -8x + -4x2

OpenStudy (anonymous):

wait...how would you get -26?